# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ 3.4 ADI-based Non-intrusive Implementation of SW-BT
$\hat{A}_{\text{SW}}$ $=W_{\text{SW}}^{*}A_{\text{SW}}\hat{V}=\hat{A}-\hat{B}D^{-1}\hat{C}$,
@@
Theorem 3.7.
Let $G(s)$ be a square stable minimum phase transfer function. Define $\tilde{L}_{w}^{*}=D^{-T}(L_{w}^{T}-\hat{B}^{*})=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,q}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,q}^{*})$. Assume that:
@@
The proof of 1 and 2 is similar to that of Theorem 3.4 and hence omitted for brevity.
@@
$\big(S_{w}-L_{w}\zeta-\hat{W}^{*}BD^{-1}\zeta\big)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}D^{-T}L_{w}^{T}=0.$
@@
Similar to DD-ADI-BT [2], the ADI-based nonintrusive implementation of SW-BT can be derived directly using the results of Theorem 3.7. The pseudo-code for this implementation, named “Data-driven SW-BT (DD-SW-BT)”, is provided in Algorithm 4.
ADI shifts for approximating $P$: $(-\sigma_{1},\cdots,-\sigma_{v})$; ADI shifts for approximating $Q_{\text{SW}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 4 $D$-SW-BT
[…]
2: Set $S_{v}$, $L_{v}$, $S_{w}$, and $L_{w}$ as in (6).
3: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
5: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
7: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
8: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
3.5 ADI/RADI-based Non-intrusive Implementation of BST
@@
The RADI-based approximation of $\tilde{Q}_{\text{S}}$ can be obtained using Theorem 3.2 as follows. The projection matrix $W_{\text{S}}$ is computed by solving the following Sylvester equation:
@@
$\hat{A}_{\text{S}}=W_{\text{\text{S}}}^{*}\tilde{A}_{\text{S}}V_{\text{S}},\quad\hat{B}_{\text{S}}=W_{\text{\text{S}}}^{*}\tilde{B}_{\text{S}},\quad\hat{C}_{\text{S}}=C_{\text{S}}V_{\text{S}},$
where $W_{\text{\text{S}}}^{*}V_{\text{S}}=I$.
@@
$=\begin{bmatrix}R_{\text{S}}^{-\frac{1}{2}}\zeta\big(\mu_{1}I-S_{w}+L_{w}\zeta+(\hat{W}^{*}\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+\hat{W}^{*}BD^{T})R_{\text{S}}^{-1}\zeta\big)^{-1}\\ \vdots\\ R_{\text{S}}^{-\frac{1}{2}}\zeta\big(\mu_{w}I-S_{w}+L_{w}\zeta+(\hat{W}^{*}\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+\hat{W}*BD^{T})R_{\text{S}}^{-1}\zeta\big)^{-1}\end{bmatrix}\hat{W}^{*}B_{\text{X}},$
$=T_{w}^{*}\hat{W}^{*}B_{\text{S}},$
@@
$-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{S}}^{-\frac{1}{2}}L_{w}^{T}=0.$
Similar to DD-ADI-BT [2], the ADI- and RADI-based nonintrusive implementation of BST can be obtained readily. The pseudo-code for this implementation, named “Data-driven BST (DD-BST)”, is provided in Algorithm 5.
ADI shifts for approximating $P$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{S}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 5 $D$-BST

### v3

@@ 3.4 Non-intrusive ADI-based Approximation of SWBT
$\hat{A}_{\text{SW}}$ $=W_{\text{SW}}^{*}A_{\text{SW}}V_{\text{SW}}=\hat{A}-\hat{B}D^{-1}\hat{C}$,
@@
Theorem 7.
Let $G(s)$ be a square stable minimum phase transfer function. Define $\tilde{L}_{w}^{*}=D^{-T}(L_{w}^{T}-\hat{B}^{*})=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,w}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,w}^{*})$. Assume that:
@@
The proof of 1 and 2 is similar to that of Theorem 4 and hence omitted for brevity.
@@
$\big(S_{w}-L_{w}\zeta-\hat{W}^{*}BD^{-1}\zeta\big)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}D^{-T}L_{w}^{T}=0.$ (76)
@@
Similar to NI-ADI-BT [61], the ADI-based non-intrusive approximation of SWBT (NI-ADI-SWBT) can be directly obtained using the results of Theorem 7. Specifically, by computing Cholesky-like factorizations $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining the low-rank factors $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}T_{w}\hat{L}_{q}$, the approximations admit the factorized representations $P\approx\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $Q_{\mathrm{SW}}\approx\tilde{L}_{q}\tilde{L}_{q}^{*}$. Substituting the exact Gramian factors $L_{p},L_{q}$ in BSA with these approximations $\tilde{L}_{p},\tilde{L}_{q}$ yields NI-ADI-SWBT, which can be implemented using transfer function samples in the right half of the $s$-plane.
3.5 Non-intrusive ADI-based Approximation of BST
@@
The RADI-based approximation of $\tilde{Q}_{\text{S}}$ can be obtained using Theorem 2 as follows. The projection matrix $W_{\text{S}}$ is computed by solving the following Sylvester equation:
@@
$\hat{A}_{\text{S}}=W_{\text{S}}^{*}\tilde{A}_{\text{S}}V_{\text{S}},\quad\hat{B}_{\text{S}}=W_{\text{S}}^{*}\tilde{B}_{\text{S}},\quad\hat{C}_{\text{S}}=C_{\text{S}}V_{\text{S}},$
where $W_{\text{S}}^{*}V_{\text{S}}=I$.
@@
$=\begin{bmatrix}R_{\text{S}}^{-\frac{1}{2}}\zeta\big(\mu_{1}I-S_{w}+L_{w}\zeta+(\hat{W}^{*}\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+\hat{W}^{*}BD^{T})R_{\text{S}}^{-1}\zeta\big)^{-1}\\ \vdots\\ R_{\text{S}}^{-\frac{1}{2}}\zeta\big(\mu_{w}I-S_{w}+L_{w}\zeta+(\hat{W}^{*}\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+\hat{W}^{*}BD^{T})R_{\text{S}}^{-1}\zeta\big)^{-1}\end{bmatrix}\hat{W}^{*}\tilde{B}_{\text{S}},$
$=T_{w}^{*}\hat{W}^{*}\tilde{B}_{\text{S}},$
@@
$-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{S}}^{-\frac{1}{2}}L_{w}^{T}=0.$ (81)
Similar to NI-ADI-BT [61], the ADI-based non-intrusive approximation of BST (NI-ADI-BST) can be obtained readily. Specifically, by computing Cholesky-like factorizations $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining the low-rank factors $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}T_{w}\hat{L}_{q}$, the approximations admit the factorized representations $P\approx\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $Q_{\mathrm{S}}\approx\tilde{L}_{q}\tilde{L}_{q}^{*}$. Substituting the exact Gramian factors $L_{p},L_{q}$ in BSA with these approximations $\tilde{L}_{p},\tilde{L}_{q}$ yields NI-ADI-BST, which can be implemented using transfer function samples in the right half of the $s$-plane.
3.6 Numerical Issues and Their Solutions
The ADI-based non-intrusive approximations presented in this section involve the inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$, which can cause numerical problems as the number of samples used for implementation increases. Specifically, as the number of samples increases, the matrices $Q_{v}$ and $P_{w}$ start to lose numerical rank. Note that these matrices do not appear explicitly in the original low-rank ADI algorithms. This section addresses this issue by providing an appropriate choice of interpolation points that guarantee the invertibility of these matrices.
Proposition 8 establishes that when $\sigma_{i}=\epsilon+j\omega_{i}$, the free parameter $\zeta$ in the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ can be set to enforce a modal structure on $\hat{A}=S_{v}-\zeta L_{v}$ such that it has poles at $\epsilon+j\omega_{i}$. The motivation for enforcing this structure is to benefit from the interesting properties of modal state-space realizations with lightly damped poles, wherein the modes are effectively decoupled from each other leading to block-diagonally dominant solutions to several matrix equations appearing in MOR and control design; see [25, 13, 14] for details.
Proposition 8.
Let $\epsilon>0$ and let distinct frequencies $\omega_{1},\dots,\omega_{v}\in\mathbb{R}$ be given such that $\omega_{i}\neq 0$ for all $i=1,\dots,v$. Define the minimum frequency magnitude as $\omega_{\min}:=\min_{i}|\omega_{i}|>0$. Consider the matrices
$S_{v}$ $=\operatorname{diag}(\epsilon+j\omega_{1},\dots,\epsilon+j\omega_{v})\otimes I_{m},$ $L_{v}$ $=\mathbf{1}_{v}^{T}\otimes I_{m},$
$\zeta$ $=(2\epsilon\mathbf{1}_{v})\otimes I_{m},$ $\hat{A}_{d}$ $=\operatorname{diag}(-\epsilon+j\omega_{1},\dots,-\epsilon+j\omega_{v})\otimes I_{m},$
where $\mathbf{1}_{v}\in\mathbb{R}^{v}$ is the column vector of ones. Let us decompose $\hat{A}=S_{v}-\zeta L_{v}$ into its block-diagonal part $\hat{A}_{d}$ and off-diagonal remainder $\hat{E}$ such that $\hat{A}=\hat{A}_{d}+\hat{E}$. Then, the following properties hold:
  1. 1.
    For any $v\geq 1$, the spectral norm of the off-diagonal approximation error is exactly
    $\|\hat{A}-\hat{A}_{d}\|_{2}=2\epsilon(v-1).$ (82)
  2. 2.
    For any $v\geq 2$, $\hat{A}$ is strictly block diagonally dominant with respect to the spectral norm provided that
    $\epsilon<\frac{\omega_{\min}}{\sqrt{4(v-1)^{2}-1}}.$ (83)
  3. 3.
    For any desired error tolerance $\delta>0$ and $v\geq 2$, choosing
    $\epsilon\leq\min\left(\frac{\delta}{2(v-1)},\;\frac{\omega_{\min}}{\sqrt{4(v-1)^{2}-1}}\right)$ (84)
    guarantees that $\|\hat{A}-\hat{A}_{d}\|_{2}\leq\delta$ and that $\hat{A}$ is strictly block diagonally dominant.
Proof.
The proof is given in Appendix A. ∎
The next proposition establishes the block diagonal dominance of $Q_{v}$ in NI-ADI-LQGBT, which solves the Lyapunov equation (30).
Proposition 9.
Let Theorem 1 and Proposition 8 hold. Define the matrix $M:=L_{v}^{*}L_{v}+\hat{C}^{*}\hat{C}$ and denote its $(i,j)$-th block entry by $M_{ij}\in\mathbb{C}^{m\times m}$. Then:
  1. 1.
    The block entries $Q_{v}]_{ij}$ of $Q_{v}$ are given by
    $Q_{v}]_{ij}=\frac{M_{ij}}{2\epsilon+j(\omega_{j}-\omega_{i})}.$ (85)
  2. 2.
    Since $L_{v}^{*}L_{v}$ has identity blocks on the diagonal, $\|M_{ii}\|_{2}\geq 1$ for all $i$. There exists an $\epsilon^{*}>0$ such that for all $0<\epsilon<\epsilon^{*}$, $Q_{v}$ is strictly block diagonally dominant with respect to the spectral norm. Specifically, dominance holds if
    $\epsilon<\min_{i}\frac{\|M_{ii}\|_{2}}{2\sum_{k\neq i}\frac{\|M_{ik}\|_{2}}{|\omega_{k}-\omega_{i}|}}.$ (86)
  3. 3.
    As $\epsilon\to 0$, the off-diagonal blocks converge to finite values determined by the frequency separation:
    $\lim_{\epsilon\to 0}[Q_{v}]_{ij}=\frac{M_{ij}}{j(\omega_{j}-\omega_{i})},\quad\text{for }i\neq j.$ (87)
    Conversely, the diagonal blocks diverge as $\|[Q_{v}]_{ii}\|_{2}\sim O(\epsilon^{-1})$.
Proof.
The proof is given in Appendix B. ∎
The next proposition establishes the block-diagonal dominance of the matrix $T_{v}$ appearing in NI-ADI-PRBT and NI-ADI-BRBT.
Proposition 10.
Let Proposition 8 hold. Let $\mathcal{P}\in\mathbb{C}^{m\times vm}$ and $\mathcal{Q}\in\mathbb{C}^{m\times m}$ be matrices independent of $\epsilon$. Decompose $\mathcal{P}$ into block columns $\mathcal{P}=[\mathcal{P}_{1},\dots,\mathcal{P}_{v}]$ with $\mathcal{P}_{k}\in\mathbb{C}^{m\times m}$. Assume that $(I_{m}+\mathcal{P}_{i})$ is invertible for all $i=1,\dots,v$ and that $\mathcal{Q}\neq 0$. Let $T_{v}\in\mathbb{C}^{vm\times vm}$ be the unique solution to the Sylvester equation:
$(S_{v}-\zeta L_{v}-\zeta\mathcal{P})T_{v}-T_{v}S_{v}+\zeta\mathcal{Q}L_{v}=0.$ (88)
Denote the $(i,j)$-th block of $T_{v}$ by $T_{v}]_{ij}\in\mathbb{C}^{m\times m}$. Then:
  1. 1.
    As $\epsilon\to 0$, the diagonal blocks converge to a finite limit, while the off-diagonal blocks vanish linearly with $\epsilon$:
    $\lim_{\epsilon\to 0}[T_{v}]_{ii}$ $=(I_{m}+\mathcal{P}_{i})^{-1}\mathcal{Q},$ (89)
    $\|[T_{v}]_{ij}\|_{2}$ $=O(\epsilon),\quad\text{for }i\neq j.$ (90)
  2. 2.
    There exists an $\epsilon^{*}>0$ such that for all $0<\epsilon<\epsilon^{*}$, $T_{v}$ is strictly block diagonally dominant with respect to the spectral norm. A sufficient condition for dominance is
    $\epsilon<\min_{i}\frac{\|(I_{m}+\mathcal{P}_{i})^{-1}\mathcal{Q}\|_{2}}{2\|\mathcal{Q}\|_{2}\sum_{k\neq i}\frac{1}{|\omega_{k}-\omega_{i}|}}.$ (91)
  3. 3.
    The error between $T_{v}$ and its block-diagonal part $\operatorname{blkdiag}([T_{v}]_{11},\dots,[T_{v}]_{vv})$ satisfies
    $\|T_{v}-\operatorname{blkdiag}(T_{v})\|_{2}\leq K\epsilon,$ (92)
    for some constant $K>0$ independent of $\epsilon$.
Proof.
The proof is given in Appendix C. ∎
Let $\sigma_{i}=\epsilon+j\omega_{i}$ and $\mu_{i}=\epsilon+j\nu_{i}$, where $\epsilon$ is determined by Propositions 8, 9, and 10 to guarantee block diagonal dominance. This selection of interpolation points effectively resolves the invertibility challenges in the proposed ADI-based non-intrusive implementations, ensuring that the matrices $Q_{v}$, $P_{w}$, $T_{v}$, and $T_{w}$ remain invertible.
If the number of samples of the transfer function to be processed is significantly high such that the Lyapunov and Sylvester equations involved in the proposed non-intrusive implementations become computationally expensive to solve, then one can exploit the block diagonal dominance properties and use the block diagonal approximations instead. As shown in [25, 13, 14], the decoupling of modes in lightly damped modal state-space realizations makes the block diagonal approximations quite effective. By exploiting the block diagonal dominance properties, the non-intrusive ADI-based implementations presented in this section can be implemented as follows.
For NI-ADI-LQGBT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{m}+H^{*}(\epsilon+j\omega_{1})H(\epsilon+j\omega_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\epsilon[I_{m}+H^{*}(\epsilon+j\omega_{v})H(\epsilon+j\omega_{v})]^{-1}\big),$ (93)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{p}+H(\epsilon+j\nu_{1})H^{*}(\epsilon+j\nu_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\epsilon[I_{p}+H(\epsilon+j\nu_{w})H^{*}(\epsilon+j\nu_{w})]^{-1}\big).$ (94)
For NI-ADI-$\mathcal{H}_{\infty}$BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{m}+(1-\gamma^{-2})H^{*}(\epsilon+j\omega_{1})H(\epsilon+j\omega_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\epsilon[I_{m}+(1-\gamma^{-2})H^{*}(\epsilon+j\omega_{v})H(\epsilon+j\omega_{v})]^{-1}\big),$ (95)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{p}+(1-\gamma^{-2})H(\epsilon+j\nu_{1})H^{*}(\epsilon+j\nu_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\epsilon[I_{p}+(1-\gamma^{-2})H(\epsilon+j\nu_{w})H^{*}(\epsilon+j\nu_{w})]^{-1}\big).$ (96)
For NI-ADI-PRBT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{m}-t_{v,1}^{*}H^{*}(\epsilon+j\omega_{1})R_{\text{PR}}^{-1}H(\epsilon+j\omega_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\epsilon[I_{m}-t_{v,v}^{*}H^{*}(\epsilon+j\omega_{v})R_{\text{PR}}^{-1}H(\epsilon+j\omega_{v})t_{v,v}]^{-1}\big),$ (97)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{p}-t_{w,1}^{*}H(\epsilon+j\nu_{1})R_{\text{PR}}^{-1}H^{*}(\epsilon+j\nu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\epsilon[I_{p}-t_{w,w}^{*}H(\epsilon+j\nu_{w})R_{\text{PR}}^{-1}H^{*}(\epsilon+j\nu_{w})t_{w,w}]^{-1}\big),$ (98)
$T_{v}$ $\approx\text{blkdiag}\big(t_{v,1},\cdots,t_{v,v}\big),$ (99)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{w,w}\big),$ (100)
where
$t_{v,i}$ $=[I_{m}+R_{\text{PR}}^{-1}H(\epsilon+j\omega_{i})]^{-1}R_{\text{PR}}^{-\frac{1}{2}},$ (101)
$t_{w,i}$ $=[I_{p}+R_{\text{PR}}^{-1}H^{*}(\epsilon+j\nu_{i})]^{-1}R_{\text{PR}}^{-\frac{1}{2}}.$ (102)
For NI-ADI-BRBT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{m}-t_{v,1}^{*}H^{*}(\epsilon+j\omega_{1})R_{p}^{-1}H(\epsilon+j\omega_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\epsilon[I_{m}-t_{v,v}^{*}H^{*}(\epsilon+j\omega_{v})R_{p}^{-1}H(\epsilon+j\omega_{v})t_{v,v}]^{-1}\big),$ (103)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{p}-t_{w,1}^{*}H(\epsilon+j\nu_{1})R_{q}^{-1}H^{*}(\epsilon+j\nu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\epsilon[I_{p}-t_{w,w}^{*}H(\epsilon+j\nu_{w})R_{q}^{-1}H^{*}(\epsilon+j\nu_{w})t_{w,w}]^{-1}\big),$ (104)
$T_{v}$ $\approx\text{blkdiag}\big(t_{v,1},\cdots,t_{v,v}\big),$ (105)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{w,w}\big),$ (106)
where
$t_{v,i}$ $=[I_{m}-D^{T}R_{p}^{-1}H(\epsilon+j\omega_{i})]^{-1}R_{b}^{\frac{1}{2}},$ (107)
$t_{w,i}$ $=[I_{p}-DR_{q}^{-1}H^{*}(\epsilon+j\nu_{i})]^{-1}R_{c}^{\frac{1}{2}}.$ (108)
For NI-ADI-SWBT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon I_{m},\cdots,2\epsilon I_{m}\big),$ (109)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon I_{p},\cdots,2\epsilon I_{p}\big),$ (110)
$T_{w}$ $\approx\text{blkdiag}\big([G(\epsilon+j\nu_{1})]^{-*},\cdots,[G(\epsilon+j\nu_{w})]^{-*}\big).$ (111)
For NI-ADI-BST:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon I_{m},\cdots,2\epsilon I_{m}\big),$ (112)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\epsilon[I_{p}-t_{w,1}^{*}\mathscr{H}(\epsilon+j\nu_{1})R_{\text{S}}^{-1}\mathscr{H}^{*}(\epsilon+j\nu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\epsilon[I_{p}-t_{w,w}^{*}\mathscr{H}(\epsilon+j\nu_{w})R_{\text{S}}^{-1}\mathscr{H}^{*}(\epsilon+j\nu_{w})t_{w,w}]^{-1}\big),$ (113)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{w,w}\big),$ (114)
where
$\mathscr{H}(\epsilon+j\nu_{i})$ $=\begin{cases}\frac{2\epsilon}{j\nu_{i}-j\omega_{i}}\big[H(\epsilon+j\omega_{i})-H(\epsilon+j\nu_{i})\big]H^{*}(\epsilon+j\omega_{i})\\ \hskip 56.9055pt+H(\epsilon+j\nu_{i})D^{T}&\text{if }\omega_{i}\neq\nu_{i}\\ 2\epsilon H^{\prime}(\epsilon+j\omega_{i})H^{*}(\epsilon+j\omega_{i})+H(\epsilon+j\nu_{i})D^{T}&\text{if }\omega_{i}=\nu_{i}\end{cases},$
$t_{w,i}$ $=\Big[R_{\text{S}}+\mathscr{H}^{*}(\epsilon+j\nu_{i})\Big]^{-1}R_{\text{S}}^{\frac{1}{2}}.$ (115)
Note that if these block diagonally dominant approximations are used, the theoretical guarantees on positive-realness, bounded-realness, and the minimum-phase property proved in Theorems 37 are lost, even if, in practice, the interpolant retains these properties.
Remark 2.